Convert the numeral to a numeral in base 10: 1100 two (the two is written smaller and is next to the 1100).
Answer by fcabanski(1391)

Check my profile and email me for one to one math tutoring online.
1100 with a subscript 2 indicates that the number is written in base-2. We normally write numbers in base ten. You should be familiar with places in base 10.
The first digit to the left of the decimal is the ones (10^0), then the next is the tens (10^1), hundreds (10^2), thousands (10^3), and it continues.
In base-19 1100 is one thousand, one hundred.
1's - 0
10's - 0
100's - 1 = 100
1000's - 1 = 1000
1000 + 100 = 1100
In base two the the first digit to the left of the decimal is the ones (2^0), then the next is the twos (2^1), then the fours (2^2) then the 8's (2^3) and it continues.
1100 in base-2 is:
1's - 0
2's - 0
4's - 1 = 4
8's - 1 = 8
That's 8+4 = 12 in base-10 notation.